# Coherence-guided Preference Disentanglement for Cross-domain Recommendations

### v1

@@ 4.3. Baseline Methods
  • BPRMF (Rendle et al., 2012): BPRMF is a famous method that presents a generic optimization criterion BPR-Opt for personalized ranking.
@@ 4.5. Overall Performance Comparison with Baseline Models (RQ1)
In Table 3 and Table 4, we provide a performance comparison of CoPD with various approaches, using metrics such as HR@10 and NDCG@10. The results indicate that the performance of CoPD significantly outperforms all competing baselines, demonstrating the effectiveness of our CoPD in solving the data sparsity problem in both domains. Based on the experimental results, we have the following observations: (1) Compared with single-domain methods, CoPD demonstrates superior performance. This indicates that our proposed global-local graph encoder framework, comprising a global graph encoder and two domain-specific local graph encoders, effectively transfers shared information across the two domains. (2) Compared with disentangle-based single-domain methods, although DGCF and DICE demonstrate outstanding performance in certain domains, outperforming several cross-domain recommendation approaches, they often yield inferior results in other domains. Our CoPD outperforms DGCF and DICE and achieves significant improvement in both domains. This indicates that CoPD produces more effective domain-shared disentanglement representations. These representations capture user-shared interest and conformity preferences and make better recommendations. (3) Compared with cross-domain methods, DisenCDR, DIDA-CDR, ETL, UniCDR and our CoPD outperform all other cross-domain methods, indicating that disentanglement-based cross-domain recommendation methods can effectively extract users’ shared information and mitigate the negative transfer problem (Zhou et al., 2023; Zhu et al., 2021b). Disentanglement-based methods achieve excellent recommendation results by transferring this shared information across different domains. (4) Compared with disentanglement-based CDR. For DisenCDR and ETL, our proposed model not only employs domain-shared user embedding to derive shared user preference but also introduces the coherence constraints to capture the domain-shared item attributes, resulting in improved recommendations. Particularly in CDR tasks with high domain similarity in Table 3, we achieve significant performance improvement compared with DisenCDR and ETL. For DIDA-CDR, which disentangles user-shared embedding into domain-shared and domain-independent embeddings, our model adopts a different disentanglement approach. CoPD disentangles user-shared embedding into conformity and interest embeddings and outperforms DIDA-CDR significantly. UniCDR employs contrastive learning to maximize mutual information between specific and shared representations, which may disrupt the original distribution of domain-specific features. CoPD apply a stop gradient mechanism when maximize mutual information between specific and shared representations, which freezes the parameters of domain-specific embeddings. This ensures that domain-specific features remain unaffected and retain their uniqueness. Therefore, CoPD achieve significant performance improvement compared with UniCDR in all CDR tasks.
@@ 5. CONCLUSION
In the future, we plan to enhance our disentanglement techniques by integrating side information, such as user reviews and item category data, to more effectively capture users’ shared preferences across domains and enhance cross-domain recommendations’ performance. Furthermore, we plan to devise specific coherence constraints for different cross-domain recommendation tasks, such as cold-start user and cold-start item recommendations. Finally, we will investigate the preference disentanglement in group-based recommendations. Motivated by the insights of this paper, we hold that users within the same group share common preferences, referred to as ’conformity‘ and ’interest‘ studied in this paper. By disentangling users’ group preferences into conformity and interest, we anticipate enhancing the performance of group-based recommendations.
@@ Acknowledgment
This work is supported by the National Natural Science Foundation of China (grant No. 62002052) and the Sichuan Science and Technology Program (Grant No. 2022YFG0189).

### v2

@@ 4.5. Overall Performance Comparison with Baseline Models (RQ1)
In Table 3 and Table 4, we provide a performance comparison of CoPD with various approaches, using metrics such as HR@10 and NDCG@10. The results indicate that the performance of CoPD significantly outperforms all competing baselines, demonstrating the effectiveness of our CoPD in solving the data sparsity problem in both domains. Based on the experimental results, we have the following observations: (1) Compared with single-domain methods, CoPD demonstrates superior performance. This indicates that our proposed global-local graph encoder framework, comprising a global graph encoder and two domain-specific local graph encoders, effectively transfers shared information across the two domains. (2) Compared with disentangle-based single-domain methods, although DGCF and DICE demonstrate outstanding performance in certain domains, outperforming several cross-domain recommendation approaches, they often yield inferior results in other domains. Our CoPD outperforms DGCF and DICE and achieves significant improvement in both domains. This indicates that CoPD produces more effective domain-shared disentanglement representations. These representations capture user-shared interest and conformity preferences and make better recommendations. (3) Compared with cross-domain methods, DisenCDR, DIDA-CDR, ETL, UniCDR and our CoPD outperform all other cross-domain methods, indicating that disentanglement-based cross-domain recommendation methods can effectively extract users’ shared information and mitigate the negative transfer problem (Zhou et al., 2023; Zhu et al., 2021c). Disentanglement-based methods achieve excellent recommendation results by transferring this shared information across different domains. (4) Compared with disentanglement-based CDR. For DisenCDR and ETL, our proposed model not only employs domain-shared user embedding to derive shared user preference but also introduces the coherence constraints to capture the domain-shared item attributes, resulting in improved recommendations. Particularly in CDR tasks with high domain similarity in Table 3, we achieve significant performance improvement compared with DisenCDR and ETL. For DIDA-CDR, which disentangles user-shared embedding into domain-shared and domain-independent embeddings, our model adopts a different disentanglement approach. CoPD disentangles user-shared embedding into conformity and interest embeddings and outperforms DIDA-CDR significantly. UniCDR employs contrastive learning to maximize mutual information between specific and shared representations, which may disrupt the original distribution of domain-specific features. CoPD apply a stop gradient mechanism when maximize mutual information between specific and shared representations, which freezes the parameters of domain-specific embeddings. This ensures that domain-specific features remain unaffected and retain their uniqueness. Therefore, CoPD achieve significant performance improvement compared with UniCDR in all CDR tasks.
@@ 5. CONCLUSION
In the future, we plan to enhance our disentanglement techniques by integrating side information, such as user reviews and item category data, to more effectively capture users’ shared preferences across domains and enhance cross-domain recommendations’ performance. Furthermore, we plan to devise specific coherence constraints for different cross-domain recommendation tasks, such as cold-start user and cold-start item recommendations. Finally, we will investigate the preference disentanglement in group-based recommendations. Motivated by the insights of this paper, we hold that users within the same group share common preferences, referred to as conformity and interest studied in this paper. By disentangling users’ group preferences into conformity and interest, we anticipate enhancing the performance of group-based recommendations.
@@ Acknowledgment
This work is supported by the National Natural Science Foundation of China (grant No. 62002052). The Shenzhen Longhua Fundamental Research Program (No. 10162A20230325B73A546). The Start-up funding provided by Charles Darwin University. The Australian Research Council partially supports this work under the streams of Future Fellowship (Grant No. FT210100624), the Discovery Project (Grant No. DP240101108), and the Linkage Project (Grant No. LP230200892).